OpenStudy (anonymous):

A motorist drives south at 28.0 m/s for 3.00 min, then turns west and travels at 25.0 m/s for 2.80 min, and finally travels northwest at 30.0 m/s for 1.00 min. For this 6.80 min trip, find the following values. a) total vector displacement ___ m (magnitude) b) __° south of west

OpenStudy (anonymous):

so far i have.. 28*3*60=5040 m 25*2.8*60=4200 m 30*1*60=1800 m now, im stuck..